2 an seven eigths divided by negative 3 an one fourth

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(21584)

2 an seven eigths divided by negative 3 an one fourth
= 23/8 div by -13/4
= 23/8 x -4/13
= -92/104
= -23/26

Answer by rfer(10417)

(23/8)/(-13/4)=
23/8*-4/13=-92/104=-23/26